Grade 12Algebra

Find
2.73 ÷ 1.3
Find
2.73 ÷ 1.3
Find
2.73 ÷ 1.3
Find
2.73 ÷ 1.3

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (273/100) ÷ (13/10)
= (273/100) × (10/13)
= (273 × 10)/ (100 × 13)
= (21 × 1)/ (10 × 1)
= 21/10
= 2.1

Thanks